OpenStudy (anonymous):

Solve for x in x-4/x=5/7

OpenStudy (anonymous):

\[\frac{ x-4 }{ x} = \frac{ 5 }{ 7}\]

geerky42 (geerky42):

Do cross-multiply; \(\dfrac{ x-4 }{ x} = \dfrac{ 5 }{ 7}~~\implies~~7(x-4) = 5x~~\implies~~7x-28 = 5x\) Can you solve for x from here?

OpenStudy (anonymous):

Yes! Thank you so much!